# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v1

@@ 1 Introduction
In this paper, we seek to quantify the increase in air pollution at open dump sites in Indonesia during 2018-2019 due to (the burning of) increased plastic waste in the wake of China’s 2018 ban. However, data scarcity presents significant obstacles to such an analysis. Indonesia, like many LMICs, has very limited ground-level (“in situ”) air quality monitoring. The country’s first reliable air quality monitors were deployed in Jakarta during our study period (Yosephine, 2016). For context: as of 2022, 47% of the world’s governments still did not collect air pollution data (OpenAQ Team, 2022). In recent years, remotely sensed data have been widely used to fill gaps in in-situ monitoring (Shin et al., 2020). In addition to the lack of air quality data, to our knowledge there is no detailed record of Indonesia’s plastic waste imports along the lines of what Unfried and Wang (2022) used to investigate domestic air quality impacts of China’s ban – and in any case, use of an official record might obscure the contribution of informal and/or covert transportation of waste, which is suggested by analysis of the Indonesian waste sector (Wang and Karasik, 2022). Therefore, we develop an analytic strategy relying on data products derived from remote sensing, including PM2.5 estimates, the identification of locations where waste has been openly dumped, and proximity to ports from which plastic waste can enter Indonesia (data sources are detailed in Section 2). To our knowledge, this is the first analysis providing empirical estimates of how a policy-driven change in plastic waste quantity – and associated changes in burning – impacts air quality, in a setting without in-situ air quality monitoring or waste quantity data.
@@ 2.4 Covariates: Meteorology, Population Density, and Fire Locations
Population density is another potential confounder in our case study. The GPW API provides an estimate of the population living within 1, 5, and 10 kilometers of each dump site. Because we observed similar patterns across the sites between these variables, we chose to adjust for population living within 1km for simplicity (it is straightforward to think of this as population density).
Lastly, we obtained the locations of active fires from VIIRS, the Visible Infrared Imaging Radiometer Suite (NASA FIRMS, 2020). The availability of this remotely sensed product (375m resolution daily) starting in 2012 influenced our selection of the study period start. A time series of the total number of active fires across Indonesia is shown in Figure S1. After obtaining the maximum footprint of each dump site as detected by GPW, we performed a spatial overlay of the active fire points. To account for imprecise detection of site boundaries as well as shape fluctuation over time, we used a 100m buffer around each site for the overlay. Throughout our analysis, we consider non-waste dump fires to be those that occurred more than 100m away from any GPW site. We use the monthly count of fires occurring outside of open dump sites in the province as a covariate in our main analysis. The purpose of this is to account for increased PM2.5 due to wildfires and agricultural burning.
@@
The VIIRS fire overlay also justifies our analysis of all dump site locations from 2012-2019, regardless of when GPW first detected them (prior to 2020). As shown in Figure S2, there were many locations at which major fires occurred prior to their date of first detection as a dump site by GPW.
@@ 3.2 Motivations for Efficient Influence Function-Based Estimation
$E[Y_{it}]=\beta_{0}+\boldsymbol{\beta}\mathbf{X}_{it}+\tau B_{t}+f(D_{i})B_{t}$
where $f(D)$ is a flexible (nonparametric) function. There are several drawbacks of this approach. To begin, even if we allow a flexible relationship between $D$ and $Y$, we are not allowing for flexible relationships between $\mathbf{X}$ and $Y$, nor interactions between the components of $\mathbf{X}$. Importantly, we are not addressing confounding between the dose $D$ and trends in the outcome – for instance, if the overall (or seasonal) trend in air pollution is affected by population density, which is static in our dataset and which we know to be correlated with $D$ (see Section 4.2). To address this, we could make a much more complicated linear regression (e.g., by using splines and more interaction terms), or use a nonparametric machine learning algorithm to model $Y$ given $D,\mathbf{X},B$. However, this outcome model might still be misspecified. This motivates the search for a doubly-robust estimator, which incorporates both a model for the outcome and a model for the propensity score (the probability of receiving treatment), which explicitly addresses confounding, and will produce results that are consistent (reliable in a probabilistic sense) if at least one of these models is correctly specified [^2] While some papers have questioned the likelihood of either the [generalized] propensity score or outcome model being correctly specified in observational studies, the most well-established alternative to doubly-robust methods for continuous exposures, which utilizes matching (Wu et al., 2024), has been shown to perform poorly when the sample size is small (Cork et al., 2025). And in our application, we only observe 356 unique values of the exposure, one per dump site..
@@ 3.4 Uncertainty Quantification
For uncertainty quantification, Hettinger et al. (2024) derive a sandwich variance estimator, but mention various limitations of that approach, including that it is challenging to incorporate uncertainty in nuisance function estimation and that extensions would be required to account for correlated data. To avoid these complications, here we use a novel variant of the nonparametric bootstrap to obtain pointwise confidence intervals for the ADT. This approach avoids making asymptotic assumptions in our finite sample setting and is designed to respect the complex spatial and temporal correlation structures in our data. Rather than discrete bootstrapping, where observations are either included in a sample or left out entirely, we use weights drawn from an Exponential distribution to allow for small but nonzero inclusion, which helps to improve the observed support of $D$ in the samples (Hettinger et al., 2024). (Note that these weights are scaled to reflect the observed sample sizes in the pre- and post-intervention time periods.) Sampling one weight per location (dump site) rather than per individual observation accounts for temporal correlation across repeated observations. An important extension we propose, motivated by our case study, is to adapt the weights to account for spatial correlation in addition to temporal.
@@ 3.4.1 Spatial Weighted Bootstrap
As shown in Figures S3 and S4, the GPW dump sites are not evenly distributed over the land, nor are the number of fires detected by VIIRS evenly distributed across sites (see our informal post-hoc analysis – the number of fires at each dump site is not a variable included in our main analysis). Especially near Jakarta (the capitol), there are some clusters of sites at which far more fires have been detected compared to sites in the rest of the country. To account for residual spatial dependence in this setting, we implement spatially-correlated sampling of the bootstrap weights.
@@ 3.5 Case Study Data Application
Lastly, while dump site size is not directly accounted for in our main analyses, we conducted several sensitivity analyses to assess how this factor might influence our results, as we thought it could be an important predictor of air pollution heterogeneity across open dump sites. As mentioned in Section 2, Global Plastic Watch estimates the area (footprint) of each dump site. In reality, the dump sites are continuously fluctuating in size, but the measurements at any given time point are less reliable than their average over time; there are also frequent gaps in the remotely sensed detection time series. Therefore, we took the area at each site to be the average across all measurements.
We did not adjust for GPW site size as a covariate in the main analysis because the area of the dump sites as measured after 1/1/18 is technically a post-treatment variable which can be affected by treatment, referred to as a “descendant” of treatment (note that only 68% of the dump sites have at least one area estimate prior to 2018, so we would have to exclude many sites, potentially biasing our causal effect curve, in order to include only pre-2018 area as a covariate). Descendants of treatment should not be conditioned on in causal inference analyses (Hernán and Robins, 2020). However, we ran a sensitivity analysis to investigate the impact of including site size (averaged across 2012-2019) as a covariate. For simplicity, this sensitivity analysis uses the same nuisance function hyperparameters and spatial range (for the bootstrap weights) as the main analysis.
@@ 5 Discussion
Third, while remotely sensed data products enable application of quasi-experimental methods in data-scarce settings, we must be conscious of their potential for measurement error / misclassification. The seemingly noisiest data product we use is the GPW dump site classifications. Because all the GPW sites detected in Indonesia were manually validated (Kruse et al., 2023), there should be no false positive classifications. However, false negative classifications are possible. This could affect our analysis in two ways. On the one hand, because we include the number of active fires in each province not overlapping with dump sites as a covariate in $\mathbf{X}$, any fires mistakenly classified as non-dump fires will reduce our estimated policy effect. On the other hand, if there is spillover of PM2.5 from false negative dump sites to nearby GPW-detected dump sites, then this could increase our estimated per-site policy effect.

### v2

@@ 1 Introduction
In this paper, we seek to quantify the increase in air pollution at open dump sites in Indonesia during 2018-2019 due to (the burning of) increased plastic waste in the wake of China’s 2018 ban. However, data scarcity presents significant obstacles to such an analysis. Indonesia, like many LMICs, has very limited ground-level (“in situ”) air quality monitoring. The country’s first reliable air quality monitors were deployed in Jakarta during our study period (Yosephine, 2016). For context: as of 2020, 51% of the world’s governments did not produce air pollution data (OpenAQ Team, 2020) [^2] Note that OpenAQ has reported updates to this statistic in 2022 and 2024.. In recent years, remotely sensed data have been widely used to fill gaps in in-situ monitoring (Shin et al., 2020). In addition to the lack of air quality data, to our knowledge there is no detailed record of Indonesia’s plastic waste imports along the lines of what Unfried and Wang (2022) used to investigate domestic air quality impacts of China’s ban – and in any case, use of an official record might obscure the contribution of informal and/or covert transportation of waste, which is suggested by analysis of the Indonesian waste sector (Wang and Karasik, 2022). Therefore, we develop an analytic strategy relying on data products derived from remote sensing, including PM2.5 estimates, the identification of locations where waste has been openly dumped, and proximity to ports from which plastic waste can enter Indonesia (data sources are detailed in Section 2). To our knowledge, this is the first analysis providing empirical estimates of how a policy-driven change in plastic waste quantity – and associated changes in burning – impacts air quality, in a setting without in-situ air quality monitoring or waste quantity data.
@@ 2.4 Covariates: Meteorology, Population Density, and Fire Locations
Population density is another potential confounder in our case study. The GPW API provides an estimate of the population living within 1, 5, and 10 kilometers of each dump site, sourced from WorldPop (Tatem, 2017). Because we observed similar patterns across the sites between these variables, we chose to adjust for population living within 1km for simplicity (it is straightforward to think of this as population density).
Lastly, we obtained the locations of active fires from VIIRS, the Visible Infrared Imaging Radiometer Suite (NASA FIRMS, 2020). The availability of this remotely sensed product (375m resolution daily) starting in 2012 influenced our selection of the study period start. A time series of the total number of active fires across Indonesia is shown in Figure S1. After obtaining the maximum footprint of each dump site as detected by GPW, we performed a spatial overlay of the active fire points. To account for imprecise detection of site boundaries as well as shape fluctuation over time, we used a 100m buffer around each site for the overlay. Throughout our analysis, we consider non-waste dump fires to be those that occurred more than 100m away from any GPW site. We use the monthly count of fires occurring outside of open dump sites in the province as a covariate in our main analysis. The purpose of this is to adjust for potential changes over time in wildfire activity and agricultural burning practices, which could also contribute to changes in PM2.5 during the study period.
@@
The VIIRS fire overlay also justifies our analysis of all dump site locations from 2012-2019, regardless of when GPW first detected them (prior to 2020). As shown in Figure S2, there were many locations at which major fires occurred prior to their date of first detection as a dump site by GPW. Having observed fires at a site before it was classified as a dump site by GPW increases the likelihood that the site was being used as a dump site prior to the GPW designation.
@@ 3.2 Motivations for Efficient Influence Function-Based Estimation
$E[Y_{it}]=\beta_{0}+\boldsymbol{\beta}\mathbf{X}_{it}+\tau B_{t}+f_{\phi}(D_{i})B_{t}$
where $f_{\phi}(D)$ is a flexible (nonparametric) function, with parameters $\phi$. There are several drawbacks of this approach. To begin, even if we allow a flexible relationship between $D$ and $Y$, we are not allowing for flexible relationships between $\mathbf{X}$ and $Y$, nor interactions between the components of $\mathbf{X}$. Importantly, we are not addressing confounding between the dose $D$ and trends in the outcome – for instance, if the overall (or seasonal) trend in air pollution is affected by population density, which is static in our dataset and which we know to be correlated with $D$ (see Section 4.2). To address this, we could make a much more complicated linear regression (e.g., by using splines and more interaction terms), or use a nonparametric machine learning algorithm to model $Y$ given $D,\mathbf{X},B$. However, this outcome model might still be misspecified. This motivates the search for a doubly-robust estimator, which incorporates both a model for the outcome and a model for the propensity score (the probability of receiving treatment), which explicitly addresses confounding, and will produce results that are consistent (reliable in a probabilistic sense) if at least one of these models is correctly specified [^3] While some papers have questioned the likelihood of either the [generalized] propensity score or outcome model being correctly specified in observational studies, the most well-established alternative to doubly-robust methods for continuous exposures, which utilizes matching (Wu et al., 2024), has been shown to perform poorly when the sample size is small (Cork et al., 2025). And in our application, we only observe 356 unique values of the exposure, one per dump site..
@@ 3.4 Uncertainty Quantification
For uncertainty quantification, Hettinger et al. (2025) derive a sandwich variance estimator, but mention various limitations of that approach, including that it relies on a normal approximation, that incorporating uncertainty in nuisance function estimation is challenging, and that extensions would be required to account for correlated data. To avoid these complications, here we use a novel variant of the nonparametric bootstrap to obtain pointwise confidence intervals for the ADT. This approach avoids making asymptotic assumptions in our finite sample setting and is designed to respect the complex spatial and temporal correlation structures in our data. Rather than discrete bootstrapping, where observations are either included in a sample or left out entirely, we use weights drawn from an Exponential distribution to allow for small but nonzero inclusion, which helps to improve the observed support of $D$ in the samples (Hettinger et al., 2025). (Note that these weights are scaled to reflect the observed sample sizes in the pre- and post-intervention time periods.) Sampling one weight per location (dump site) rather than per individual observation accounts for temporal correlation across repeated observations. An important extension we propose, motivated by our case study, is to adapt the weights to account for spatial correlation in addition to temporal.
@@ 3.4.1 Spatial Weighted Bootstrap
Because our set of covariates likely does not explain all spatial correlation in the data, we need to account for residual spatial dependence in our uncertainty quantification. For instance, note that there are some clusters of dump sites near Jakarta (shown in Figure S4) which we would be hard-pressed to call independent of one another, even conditional on our covariate set. Therefore, in our weighted bootstrap procedure we would like a site’s weight to be larger (or smaller) if other sites nearby have larger (or smaller) weights. We achieve this through spatially-correlated sampling of the bootstrap weights.
@@ 3.4.2 Estimating the Residual Spatial Correlation Function
3.4.3 Simulations to Investigate the Spatial Weighted Bootstrap
A pseudo-simulation study was conducted to evaluate the properties of the proposed Spatial Weighted Bootstrap confidence intervals. In Appendix C, we provide full details on the data generating process, the calculation of performance metrics, and the results of this simulation study. Crucially, these simulations illustrate that our Spatial Weighted Bootstrap improves coverage of the confidence intervals relative to a Non-spatial Weighted Bootstrap, across a variety of scenarios.
@@ 3.5 Case Study Data Application
Lastly, while dump site size is not directly accounted for in our main analyses, we conducted a sensitivity analysis to assess how this factor might influence our results, as we thought it could be an important predictor of air pollution heterogeneity across open dump sites. As mentioned in Section 2, Global Plastic Watch estimates the area (footprint) of each dump site. In reality, the dump sites are continuously fluctuating in size, but the measurements at any given time point are less reliable than their average over time; there are also frequent gaps in the remotely sensed detection time series. Therefore, we took the area at each site to be the average across all measurements.
We did not adjust for GPW site size as a covariate in the main analysis because the area of the dump sites as measured after 1/1/18 is technically a post-treatment variable which can be affected by treatment, referred to as a “descendant” of treatment (note that only 68% of the dump sites have at least one area estimate prior to 2018, so we would have to exclude many sites, potentially biasing our causal effect curve, in order to include only pre-2018 area as a covariate). Descendants of treatment should not be conditioned on in causal inference analyses (Hernán and Robins, 2020). We ran a sensitivity analysis to investigate the impact of including site size (averaged across 2012-2019) as a covariate. For simplicity, this sensitivity analysis uses the same nuisance function hyperparameters and spatial range (for the bootstrap weights) as the main analysis.
@@ 5 Discussion
Third, while remotely sensed data products enable application of quasi-experimental methods in data-scarce settings, we must be conscious of their potential for measurement error / misclassification. In this case study, we believe that the GPW dump site classifications are the data product in which errors would be most consequential for our analysis. Because all the GPW sites detected in Indonesia were manually validated (Kruse et al., 2023), there should be no false positive classifications. However, false negative classifications are possible and could affect our analysis in two ways. On the one hand, because we include the number of active fires in each province not overlapping with dump sites as a covariate in $\mathbf{X}$, any fires mistakenly classified as non-dump fires will reduce our estimated policy effect. On the other hand, if there is spillover of PM2.5 from false negative dump sites to nearby GPW-detected dump sites, then this could increase our estimated per-site policy effect.
@@ Appendix B Supplementary Figures and Tables
Appendix C Simulations Exploring the Spatial Weighted Bootstrap
In this appendix, our primary aim is to evaluate the characteristics of the confidence intervals generated by our Spatial Weighted Bootstrap procedure, relative to those generated by a Non-spatial Weighted Bootstrap. Here, calculation of the ADT curve’s bias and Mean Absolute Error (MAE) is an intermediate step to better understand the confidence intervals, as Hettinger et al. (2025) present extensive investigation of a very similar ADT estimator.
To do this evaluation, we develop simulations which utilize the real (observed) exposure $D$ and covariates $\mathbf{X}$ to simulate the outcome $Y$.
C.1 Simulating the Outcome
First, we compute the mean function for the outcome data generating process using the real exposure and covariate data and setting “true” parameter values governing the relationship between these variables and the outcome in our simulated data. Our choice of these parameter values was guided by (a) using domain knowledge to select the sign of each coefficient, (b) producing an outcome distribution similar to that in the real data, and (c) producing a plausible ADT curve, similar to that observed in our main analysis.
The mean function for each location $i$ and time $t$ is given by:
$\mu_{it}=B_{t}L_{1}(\mathbf{X}_{it},D_{it})+(1-B_{t})L_{0}(\mathbf{X}_{it})$
where
$L_{0}(\mathbf{X}_{it})$ $=-9.5+0.075(\text{Temp}_{it})+0.075(\text{Pop}_{i})+0.0001(\text{Temp}_{it})(\text{Pop}_{i})+0.15(\text{Time}_{t})+$
            $\>\>\>\>\>\>\>\>\>\>\>\>0.001(\text{Time}_{t})(\text{Pop}_{i})+1(\text{Other\_fires\_province}_{it})-18(\text{Precip}_{it})$
and
$L_{1}(\mathbf{X}_{it},D_{it})$ $=-12.5+0.075(\text{Temp}_{it})+0.075(\text{Pop}_{i})+0.0007(\text{Temp}_{it})(\text{Pop}_{i})+0.15(\text{Time}_{t})+$
            $\>\>\>\>\>\>\>\>\>\>\>\>0.001(\text{Time}_{t})(\text{Pop}_{i})+1.5(\text{Other\_fires\_province}_{it})-20(\text{Precip}_{it})+$
            $\>\>\>\>\>\>\>\>\>\>\>\>2(D_{it})\Big{[}-3(D_{it})+0.4(\text{Pop}_{i})+0.2(\text{Time}_{t})\Big{]}$
where “Temp” is temperature, “Pop” is population [density], “Time” is the fifth-root of the count of days since 01/01/2012, “Other_fires_province” is the fourth-root of the number of fires in the province not overlapping with any GPW site, and “Precip” is precipitation. For both $L_{0}$ and $L_{1}$, we truncated any negative values ($\sim$0.06%) to zero post hoc.
We calculate the “true” ADT for each exposure value $\delta$ as
$\Psi(\delta)=\frac{1}{96N}\sum_{i=1}^{N}\sum_{t=1}^{96}\ \Big{[}L_{1}(\mathbf{X}_{it},\delta)-L_{0}(\mathbf{X}_{it})\Big{]}$
This true ADT curve is shown in Figure S12.
We simulate both spatially-structured and independent error in the outcome, as follows:
$\tilde{Y}_{it}=\mu_{it}+\sigma_{\text{sp}}\epsilon_{\text{sp}}+\sigma_{\text{ind}}\epsilon_{\text{ind}}$
where $\epsilon_{\text{sp}}\sim MVN(0,\Sigma)$, where the correlation function to generate the entries of $\Sigma$ is $Corr(m;L)=exp(-m/L)$ [Exponential form] and $\epsilon_{\text{ind}}\sim N(0,1)$. The parameters $L$ (the true spatial range), $\sigma_{\text{sp}}$, and $\sigma_{\text{ind}}$ are constants, which we specify to obtain four distinct simulation scenarios:
  1. 1.
    Scenario 1: $L=100$ km, $\sigma_{\text{sp}}=2$, $\sigma_{\text{ind}}=0.5$ (the “base case”)
  2. 2.
    Scenario 2: $L=500$ km, $\sigma_{\text{sp}}=2$, $\sigma_{\text{ind}}=0.5$ (increased spatial range)
  3. 3.
    Scenario 3: $L=100$ km, $\sigma_{\text{sp}}=3$, $\sigma_{\text{ind}}=0.5$ (increased spatial sigma)
  4. 4.
    Scenario 4: $L=100$ km, $\sigma_{\text{sp}}=2$, $\sigma_{\text{ind}}=1.5$ (increased noise)
These four scenarios allow us to investigate the influence of different spatial error structures on the accuracy of our estimation and inference procedures. For each simulation scenario, we generate 100 simulated datasets (using 100 different seeds for random number generation), each having the same true parameter values and (real) exposure and covariates but different synthetic outcome variables.
C.2 Estimating the Effect Curve and Confidence Interval
When implementing our ADT estimation procedure on each simulated dataset, we correctly specify the outcome models’ mean function but estimate the propensity score models $\pi_{B}$ and $\pi_{D}$ using Ranger (the same algorithm as in our main analysis).
When applying the Spatial Weighted Bootstrap, we fit both Exponential and Power covariance functions (the latter has correlation function $Corr(m;a,z)=1-zm^{a}$) to the empirical variogram of the residuals from the LLKR stage and use whichever model has the smaller RMSE to generate the spatial weights. As noted in Section 3.4.2, the spatial structure of the LLKR residuals is different than the spatial structure of the outcome (here, $\tilde{Y}$), so traditional notions of correctly vs incorrectly specifying the spatial correlation model do not directly apply here.
Figure S12: Examples of the ADT estimate from our multiply robust estimation procedure applied to (four instances of) simulated data from our base case scenario, compared to the true ADT curve.
C.3 Evaluating the Performance of the Spatial and Non-spatial Weighted Bootstraps
We compare our ADT curve estimates from each simulated dataset to the true ADT curve by computing the pointwise bias and MAE, which are then summarized by averaging across all 100 simulation replicates within each scenario. We then calculate the pointwise coverage and width of the 95% confidence intervals computed using the Spatial Weighted Bootstrap, and compare the results with those from a Non-spatial Weighted Bootstrap.
C.4 Simulation Results and Discussion
The bias and MAE results are shown in Figures S13 and S14, respectively. The confidence interval coverage and width results are shown in Figures S15 and S16, respectively.
First, we observe that the magnitude of the bias ranges between 0.1 and 0.5, while the true ADT ranges between 0.1 and 2.6, with an average value of 1.9. From the MAE plot, we see that the estimator is less stable on the tails (resulting in larger errors), which is to be expected for methods like LLKR. Our intuition is that the asymmetry of the true ADT curve contributes to the bias being higher at higher values of port proximity.
Note that in addition to the finite sample bias of LLKR, which results in imprecise centering of the ADT’s confidence intervals (Hettinger et al., 2025), here we also do not know the correct specification of the propensity score models $\pi_{B}$ and $\pi_{D}$, instead estimating them from the data. Foreshadowing our coverage results, the simulations in Hettinger et al. (2025) indicate that misspecification of the propensity score models has a greater impact on coverage than misspecification of the outcome models.
In Figure S15, we see that neither the Spatial nor Non-spatial Weighted Bootstrap achieves 95% coverage at any value of $\delta$ (port proximity). However, the Spatial Weighted Bootstrap has consistently higher coverage than the Non-spatial method. This makes sense, as the Spatial method generates wider confidence intervals (as shown in Figure S16). Mirroring the results of the MAE plot, we see that the confidence interval coverage tends to be highest in the lower-middle range of $\delta$ (0.25-0.5) for both the Spatial and Non-spatial Weighted Bootstrap.
Now, we compare the differences in results across our four simulation scenarios. Note that both Scenario 2 (increasing $L$) and Scenario 3 (increasing $\sigma_{\text{sp}}$) generate greater similarity in the value of outcomes that are spatially proximate, through different mechanisms, relative to Scenario 1. Whereas, Scenario 4 simply increases the random noise in the outcomes.
In Figures S15 and S16, we see that inducing greater spatial signal (Scenarios 2 and 3) increases the width of the Spatial Weighted Bootstrap’s confidence intervals, resulting in higher coverage. By contrast, the Non-spatial approach responds to increasing $L$ by narrowing its confidence intervals (resulting in lower coverage) and responds to increasing $\sigma_{\text{sp}}$ by widening its confidence intervals (resulting in higher coverage). In Scenario 4, the Spatial approach actually narrows its confidence intervals, likely because increasing $\sigma_{\text{ind}}$ makes it harder to detect the spatially structured error, resulting in under-estimation of the spatial correlation.
In Figure S14, the only parameter configuration that stands out is Scenario 3 (increased $\sigma_{\text{sp}}$), which results in higher MAE (lower precision in the point estimate) across values of $\delta$. In Figure S13, none of the configurations have consistently different bias.
Figure S13: Bias of the estimated ADT compared to the true ADT in our simulations.
Figure S14: Mean Absolute Error (MAE) of the estimated ADT compared to the true ADT in our simulations.
Figure S15: Coverage of the 95% confidence intervals in our simulations. The dashed line indicates the desired 95% coverage.
Figure S16: Width of the confidence intervals in our simulations.